I have a word problem that I cannot get into the right form. If I can get it into the right form I can solve the problem. The problem is "Lisa participated in a triathlon in which she swam 3 miles, ran 5 miles, and then bicycles 10 miles. Lisa ran twice as fast as she swam, and she cycled three times as fast as she swam. If her total time for the triathlon was 1 hour and 46 minutes, then how fast did she swim?"
Can you help me to understand how to right the problem into an equation to find the answer?

Answer by rwm(914)

r*t=d
s=s
r=2s
c=3s
t=1 46/60
S+R+C=t
distances
s*S=3
r*R=5
c*C=10
rates
s=5 mph
r=10 mph
c=15 mph
times
S = 3/5 hr
R = 1/2 hr
C = 2/3 hr
swim
5*3/5=3 miles
ran
10*1/2=5 miles
cycle
15*2/3=10 miles